OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values

OpenStudy (anonymous):

just place a value into x and solve for example y = -2(0) - 3 y = -3 now if you do 1 in x y = -2(1) - 3 y = -5

OpenStudy (anonymous):

im confused...so y=-5 would be my answer??